A golf ball is dropped onto a concrete floor and strikes the floor at 5.0 m/s. It then rebounds at 5.0 m/s. calculate the change in velocity for the ball
1
Expert's answer
2018-04-24T09:40:08-0400
If we choose upwards as the positive direction, the change in velocity for the ball is
∆v=v_f-v_i=5-(-5)=10 m/s